OpenStudy (anonymous):

Medal and thank you for answering this >.< Solve for x. 12^x-2 = 3^2x+1

OpenStudy (anonymous):

let me make sure you typed that correctly or I understand it correctly \[12^x-2 = 3^{2x}+1 \]

OpenStudy (anonymous):

\[12^{x-2}=3^{2x+1}\]

OpenStudy (anonymous):

the second one >.<

OpenStudy (anonymous):

@celestialdictator

OpenStudy (anonymous):

do you know logarithms?

OpenStudy (anonymous):

@celestialdictator sorry I was eating >.< and I do but this problem is hard for me

OpenStudy (anonymous):

@ChristopherToni please help!

OpenStudy (anonymous):

try to study again

OpenStudy (anonymous):

I do know the basics. I'll reread it and hopefully find a way to solve it. Thanks anyway.

OpenStudy (anonymous):

okay !

OpenStudy (anonymous):

@PFEH.1999 Please help. I wonder if you can give a little more in depth explanation. Thank you in advance.

OpenStudy (anonymous):

@ramit.dour @neoh147 please help >.<

OpenStudy (anonymous):

wait, calculating

OpenStudy (anonymous):

\[12^{x-2}=3^{2x+1}\] \[4^{x-2}=3^{2x+1} \div 3^{x-2}\]\[4^{x-2}=3^{x-1}\]

OpenStudy (anonymous):

wait for next part of solution

OpenStudy (anonymous):

Thank you both of you so, so much. <33 Much appreciated!

OpenStudy (anonymous):

12^(x-2) = 3 ^(2x+1) (4*3)^(x-2)=3 ^(2x+1) 4^(x-2)*3^(x-2)=3 ^(2x+1) (4^x/4^2)*(3^x/3^2)=3^2x*3 (4^x/4^2)*(3^x/3^2)=(3^x)^2*3 divided both side with 3^x (4^x/4^2)*(1/3^2)=3^x*3 rearrange 4^x/3^x=3^3*4^2 (4/3)^x=3^3*4^2 x= log (4/3) 3^3*4^2 4/3 is the base change base x= (log 3^3*4^2) / log (4/4) use calculator....

OpenStudy (anonymous):

sorry last line typo error.... x= (log 3^3*4^2) / log (4/3) after press calculator, x=21.09

OpenStudy (anonymous):

now take log both sides \[\ln 4^{x-2}=\ln 3^{2x+1-(x-2)}\] (x-2)ln4=(x-1)ln3 x=(2ln4-ln3)/(-ln3+ln4)

OpenStudy (anonymous):

x=(4ln(2)-ln(3))/(2ln(2)-ln(3)) = 5.818

OpenStudy (anonymous):

Thank you so much for your help!! I'll rework the problem with your assistance to understand this. You guys are so awesome! 3x medals

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=solve+12%5E%28x-2%29+%3D+3%5E%282x%2B1%29 ans from wolfram alpha show that is 21.09...

OpenStudy (anonymous):

@neoh147 I see, thank you so much! <3

OpenStudy (anonymous):

you are welcome @study100